patrick drove 5 hours to attend a meeting. on the return trip, his speed was 10 miles per hour less and took 6 hours. what was his speed each way? how far did he drive to the meeting? Answer by mananth(16949) 
 patrick drove 5 hours to attend a meeting. on the return trip, his speed was 10 miles per hour less and took 6 hours. what was his speed each way? how far did he drive to the meeting?
Going out speed xmph
returning speed =10mph slower
returning speed x-10
returning time = 6 hours
going out time= 5hours
Distance = same
5x=6(x-10)
5x=6x-60
5x-6x=-60
-1x=-60
/-1
x=60mph